OpenStudy (anonymous):

[6.02] For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12

OpenStudy (diyadiya):

-x=-4+2y Multiply both the sides by -1 to get x x=4-2y

OpenStudy (ash2326):

@LoveYou*69 Did you understand it?

OpenStudy (diyadiya):

Now substitute it in second equation 3(4-2y) + y = 12

OpenStudy (anonymous):

Okay! Thanks! and yea ash i knew the answer i was just making sure it was right lol :)

OpenStudy (anonymous):

so is the answer -2y+4